An average worker in Brazil can produce an ounce of soybeans in 20 minutes and an ounce of coffee in 60 minutes, while an average worker in Peru can produce an ounce of soybeans in 80 minutes and an ounce of coffee in 120 minutes.

  1. Of the following four terms of trade, which is both (i) beneficial to both Brazil and Peru and (ii) most desirable from the perspective of Peru?

    1 lb. of coffee for 3.25 lbs. of soybeans

    1 lb. of coffee for 2.5 lbs. of soybeans

    1 lb. of coffee for 1.75 lbs. of soybeans

    1 lb. of coffee for 1 lb. of soybeans
Answer #1

The opportunity cost of coffee in Brazil is 20/60 = 0.33 lbs of soybean. i.e. to produce a coffee in Brazil we have to give up 0.33 lbs of Soybean. The opportunity cost of soybean is 3 lbs of Coffee per soybean.

The opportunity cost of coffee in Peru = 80/120 = 0.66 lbs of soybean. The opportunity cost of soybean is 120 /80 = 1.5.

Peru will gain from trade when they are exporting soybean. Peru can produce 1.5 lbs of soybean for every pound of coffee.they will benefit when they are giving less than 1.5 pounds of soybean for a pound of coffee. Brazil will benefit when they are getting more than 0.33 lbs of soybean for a pound of coffee.

Beneficial for both and most beneficial for Peru is  = option D.